OpenStudy (anonymous):

what is the solution set to the following system? x+y=5 x^2+y^2=25

OpenStudy (anonymous):

(0,-5) (-5,0) (0,5) (-5,0) (0,-5) (5,0) (0,5) (5,0)

OpenStudy (mathmale):

What have YOU thought of doing so far? As for me, I'd use SUBSTITUTION to solve this problem. Soplve the first equation for either x or y in terms of the other variable. Substitute the resulting expression into the 2nd equation.

OpenStudy (anonymous):

alright so x-y=5 would be y=5-x I then would plug this into x^2+y^2=25 right? so x^2 + (5-x)^2=25 yeah?

OpenStudy (anonymous):

x^2+5-x^2=25 ....?

OpenStudy (mathmale):

Nice work!\[x^2 + (5-x)^2=25\rightarrow x ^{2}+25-10x+x ^{2}=25.\]Solve for x. Once you have x, calculate y. Be certain to check all possibilities, positive and negative.

OpenStudy (mathmale):

Be careful: x^2 + (5-x)^2=25 does not result in x^2+5-x^2 = 25.

OpenStudy (anonymous):

ok so would this come out to be 2x^2 +25 -10x after adding like terms? that cant be right!